[Brahms-l] Shift report 20040314 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 14 2004 - 00:09:07 EST
------------------------------------------------------------
Date: Sun Mar 14 00:09:06 2004, Shift: 16:00-24:00
Supervisor: CEJ, with: 
Log book#15 pages 87 - 87

Shift Summary:
------------------------------------------------------------
1) Beam untill 17:30. Again from 18:15 to 22:15. 
   Now they are fixing front-end coputer problems. 

2) Pedestal runs 10981 (TPC) and 10982.
   Finished FS 10deg A306 and MRS 90 deg B350 settings.
   Took FS zero field runs and moved FS to 8deg.
   Now running full field (FS:A, MRS:B).

3) D1 went into stby-fault several times after it
   was set to full field.
   The audio alarm on the pii4 does not work! 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 14 00:09:47 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 14 2004 - 00:10:07 EST